# Elucidated Rolling Diffusion Models for Probabilistic Weather Forecasting

### v1

@@ Abstract
Diffusion models are a powerful tool for probabilistic forecasting, yet most applications in high-dimensional chaotic systems predict future snapshots one-by-one. This common approach struggles to model complex temporal dependencies and fails to explicitly account for the progressive growth of uncertainty inherent to such systems. While rolling diffusion frameworks, which apply increasing noise to forecasts at longer lead times, have been proposed to address this, their integration with state-of-the-art, high-fidelity diffusion techniques remains a significant challenge. We tackle this problem by introducing Elucidated Rolling Diffusion Models (ERDM), the first framework to successfully unify a rolling forecast structure with the principled, performant design of Elucidated Diffusion Models (EDM). To do this, we adapt the core EDM components–its noise schedule, network preconditioning, and Heun sampler–to the rolling forecast setting. The success of this integration is driven by three key contributions: $(i)$ a novel loss weighting scheme that focuses model capacity on the mid-range forecast horizons where determinism gives way to stochasticity; $(ii)$ an efficient initialization strategy using a pre-trained EDM for the initial window; and $(iii)$ a bespoke hybrid sequence architecture for robust spatiotemporal feature extraction under progressive denoising. On 2D Navier–Stokes simulations and ERA5 global weather forecasting at $1.5^{\circ}$ resolution, ERDM consistently outperforms key diffusion-based baselines, including conditional autoregressive EDM. ERDM offers a flexible and powerful general framework for tackling diffusion-based sequence generation problems where modeling escalating uncertainty is paramount. [^1] Code is available at: https://github.com/salvaRC/erdm
@@ 1 Introduction
Probabilistic forecasting of chaotic, dynamical systems is essential for realistically assessing future states (“snapshots”) and their inherent uncertainties [2, 11]. For example, medium-range weather forecasting ($\leq 15$ days) grapples with the inherent chaotic nature of the atmosphere and high sensitivities to the initial conditions. Ensembles of numerical weather prediction (NWP) models are therefore standard practice for estimating forecast uncertainty and the likelihood of high-impact events [39, 3]. However, each ensemble run of a flagship system such as IFS ENS [13] is computationally expensive, motivating data-driven alternatives [7].
@@
Diffusion models have shown notable success in generative modeling, particularly for high-dimensional data such as images and videos [49, 12]. The Elucidated Diffusion Model (EDM) framework [23], for example, enables stable high-fidelity sample generation due to its principled design that generalizes and improves on the foundational denoising diffusion probabilistic model (DDPM) [18] and related diffusion paradigms [50, 51]. This success motivates adapting diffusion models and EDM to model time-evolving dynamical systems.
@@
Autoregressive conditional diffusion models have been successfully applied to fluid dynamics [26] and weather [42]. These approaches often use limited temporal context and predict future snapshots one by one. To explicitly model the progressively increasing uncertainty common to sequence generation tasks, Rolling Sequence Diffusion Models (RSDM) [46, 54] employ snapshot-dependent noise schedules that escalate noise for distant future snapshots. However, existing RSDM’s are derived from DDPM, rather than leveraging improved practical design choices, such as network preconditioning and second-order sampling, that were key to EDM’s success.
@@
Our work addresses these limitations by building upon the successful EDM framework [23]. We adapt and extend EDM’s core design elements—its noise schedule formulation, denoiser network parameterization, network preconditioning, loss weighting, and sampling algorithms—to the specific demands of RSDM-like sequence modeling. In developing ERDM, we identified and addressed key design choices for progressive noise schedules and temporal loss weighting within the EDM paradigm, aspects that received limited attention in prior work. Our work thus provides novel solutions for these crucial components. All in all, our key contributions are summarized as follows:
@@
Figure 1: ERDM sampling with a window, highlighted in bold purple, of size $W=4$. Top row: ERDM starts at diffusion time ${t}=0$ with snapshots, ${\bm{x}}_{1},\dots,{\bm{x}}_{W}$, corrupted by progressively larger noise levels, $\bar{\sigma}_{1}(t)<\dots<\bar{\sigma}_{W}(t)=\sigma_{\text{max}}$. Middle row: After $N=2$ joint denoising steps, the sequence reaches lower noise levels at ${t}=1$ such that $\bar{\sigma}_{1}(1)=\sigma_{\text{min}}$ and $\bar{\sigma}_{w}(1)=\bar{\sigma}_{w-1}(0)$ for $w>1$, as illustrated in the right-hand panel. The now fully denoised first snapshot, ${\bm{x}}_{1}$, is returned. Bottom row: The rest of the sequence is shifted one slot to the right, and a fresh pure-noise snapshot is appended to the new window. The cycle then repeats.
@@ 2 Related Works
Diffusion models, initially transformative in image and audio generation, are increasingly being adapted for spatiotemporal modeling. Latent-diffusion variants [53, 44] provide the computational advantages integral to state-of-the-art video generators [19, 17, 5]. Recent work refines the diffusion process itself to strengthen temporal structure and boost efficiency [47, 29]. Rolling sequence diffusion models (RSDM) advance this idea by assigning progressively larger noise levels to later timesteps, mirroring the growth of predictive uncertainty [54, 46, 24]. We embrace the same intuition but depart from these DDPM-based RSDM’s in two crucial ways. (i) Our approach is derived from the EDM framework, which generalises and improves upon DDPM. (ii) We re-examine often-overlooked design choices—noise schedule, loss weighting, and network architecture—and show that their coordination is critical for forecasting chaotic dynamics. Progressive schedules like ours form a specific instance of the completely randomized schedules proposed in Chen et al. [9], corresponding to its ‘pyramid sampling’ scheme. However, our ablations indicate that such training-time randomization degrades performance–possibly because it misaligns with the temporally increasing uncertainty inherent to the chaotic systems that we study.
@@
Initial machine learning (ML)-based attempts in this area were limited to deterministic forecasting [40, 4, 27, 35, 36, 6], often differentiated by their underlying neural network architectures [22]. More recently diffusion models have shown promise for probabilistic weather forecasting [42, 1], and other tasks such as downscaling [33, 31, 10, 52], data assimilation [20, 45, 32], and emulation [28, 48, 41]. Latent-variable models offer an alternative probabilistic route [37], and physics-ML hybrids such as NeuralGCM show promise but face scalability constraints [25]. Of particular relevance to our work is GenCast [42], which successfully applied the EDM framework, using the graph architecture from [27], for next-step probabilistic weather forecasting. Our primary baseline in this paper, EDM, corresponds to the same approach but using our experimental setup (e.g., same U-Net architecture and dataset resolution; see Section D.4 for more details). This design ensures a controlled and direct comparison between ERDM and an EDM-based next-step forecasting approach.
@@ 3 Background
In the following, we refer to clean data as ${\bm{y}}$, noisy data as ${\bm{x}}$, and we use $\bar{\bm{x}}$ to make clear when the data is corrupted according to progressively increasing noise levels. We abbreviate sequences by ${\bm{y}}_{1:W}:={\bm{y}}_{1},\dots,{\bm{y}}_{W}$. Probabilistic forecasting is concerned with predicting a future sequence $\{{\bm{y}}_{\tau}\}_{\tau=1}^{T_{\mathrm{forecast}}}$ given an initial condition ${\bm{y}}_{0}$, where the horizon $T_{\mathrm{forecast}}$ can be arbitrarily large. This requires learning the joint conditional distribution, which can be approximated with, e.g., diffusion models, by $p_{\theta}({\bm{y}}_{1:T_{\mathrm{forecast}}}|{\bm{y}}_{0})$. Directly modeling the high-dimensional joint distribution is often intractable. A common alternative is modeling a smaller window of size $W$ by learning a short-term transition, $p_{\theta}({\bm{y}}_{w+1:w+W}|{\bm{y}}_{w})$ [42, 17]. Long sequences are then generated autoregressively.
@@
Conditional diffusion models can be applied to this task by learning a denoiser network $D_{\theta}({\bm{x}}_{1:W};\sigma,{\bm{y}}_{0})$. The denoiser is trained to predict the target sequence, ${\bm{y}}_{1:W}$, given the initial condition, ${\bm{y}}_{0}$, and data ${\bm{x}}_{1:W}$ corrupted with Gaussian noise of standard deviation $\sigma\in[\sigma_{\text{min}},\sigma_{\text{max}}]$. The maximum noise level needs to be chosen such that $\sigma_{\text{max}}\gg\sigma_{\text{data}}$, where $\sigma_{\text{data}}$ is the standard deviation of the data ${\bm{y}}\sim p_{\rm{data}}$ and ${\bm{y}}\in\mathbb{R}^{D}$. The trained network can then be used jointly with an ODE or SDE solver to iteratively generate the clean data starting from pure noise, $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\sigma_{\text{max}}^{2}\mathbf{I}_{W\times D})$, and ${\bm{y}}_{0}$. See Appendix B for additional background on diffusion. Often, $W=1$ [42, 26], and the diffusion model becomes a next-step forecasting model. However, the short window size models limited temporal interactions and generating each step requires a full reverse diffusion process, incurring substantial computational cost. Standard sequence diffusion models with $W>1$ typically apply noise uniformly across the prediction window, failing to capture the increasing uncertainty inherent to future predictions.
@@
Rolling Sequence Diffusion Models (RSDMs) [46, 9] address prior limitations by explicitly incorporating progressive forecasting uncertainty. The core innovation is a progressive noise schedule with $W>1$. Instead of uniform noise, RSDMs use a monotonically increasing schedule of noise standard deviations $0\approx\bar{\sigma}_{1}^{\text{}}({t})<\bar{\sigma}_{2}^{\text{}}({t})<\dots<\bar{\sigma}_{W}^{\text{}}({t})\approx\sigma_{\text{max}}$, where ${t}$ denotes the global diffusion time. A noisy future state is thus modeled as $\bar{\bm{x}}_{w}\sim\mathcal{N}({\bm{y}}_{w},\bar{\sigma}^{2}_{w}({t})\mathbf{I}_{D})$. This structure reflects higher uncertainty for predictions further into the future. The denoiser $D_{\theta}(\bar{\bm{x}}_{1:W};\bm{\bar{\sigma}}_{1:W}(t))$ is trained analogously to a standard diffusion model to predict the clean sequence, ${\bm{y}}_{1:W}$. In our implementation, we omit direct conditioning of the denoiser on the last known clean state, ${\bm{y}}_{0}$. This is feasible because we initialize the first window using an external forecaster (itself conditioned on ${\bm{y}}_{0}$). The minimally noisy start of the window provides sufficient conditioning to predict subsequent snapshots. When requiring the RSDM to self-initialize its first window, this is not possible [46].
@@ 4 Forecasting with Elucidated Rolling Diffusion Models
Our work, ERDM, integrates the rolling diffusion concept with the Elucidated Diffusion Model (EDM) framework [23]. While RSDMs introduce the idea of progressive noise, they usually do not follow EDM’s principled design choices for preconditioning, loss weighting, and sampler design, which we hypothesize could bring benefits when incorporated into RSDMs too. ERDM aims to systematically port and adapt these EDM principles to the rolling, progressive noise setting, thereby providing a robust and theoretically grounded approach to sequential data generation.
@@
Figure 2: Noise schedule comparison for a sequence length $W=6$, yielding 6 visualized segments $[\bar{\sigma}_{w}(1),\bar{\sigma}_{w}(0)]$. The schedules differ by their curvature parameter $\rho$: one using $\rho=7$ (default EDM) and the other $\rho=-10$ (ERDM). Color gradients illustrate segment progression.
where $t_{w,t}=1-\frac{w-t}{W}$ is the local diffusion time of the snapshot. The $1-$ is important so that near-future snapshots receive minimal amounts of noise, while far-future snapshots receive high levels of noise. This ensures the correct inductive bias that distant snapshots are more uncertain than near-future ones. The continuous noise schedule is effectively divided into $W$ segments such that the noise levels at snapshot $w$ satisfy $\bar{\sigma}_{w}(t)\in[\bar{\sigma}_{w-1}(0),\bar{\sigma}_{w+1}(1)]$, where $t\in[0,1],\bar{\sigma}_{0}(0):=\sigma_{\text{min}},$ and $\bar{\sigma}_{W+1}(1):=\sigma_{\text{max}}$. Our formulation thus satisfies $\sigma_{\text{min}}=\bar{\sigma}_{1}(1)<\bar{\sigma}_{1}(0)=\bar{\sigma}_{2}(1)<\dots<\bar{\sigma}_{W}(0)=\sigma_{\text{max}}$. Two example noise schedules are illustrated in Fig. 2. ERDM is trained to see any noise schedule level within the noise segments (color gradients in Fig. 2) by randomizing $t\in U([0,1))$. During sampling, each $\bar{\sigma}_{w}({t})$ is evolved from $\bar{\sigma}_{w}(0)$ to $\bar{\sigma}_{w}(1)$. We found it important to tune the curvature parameter, $\rho$. Our proposed default value, $\rho=-10$, effectively causes snapshots to be under less noise compared to the default EDM choice, $\rho=7$, which provides more information for accurate denoising. Importantly, these noise schedule parameters $\sigma_{\text{min}},\sigma_{\text{max}},$ and $\rho$ need to be chosen at training time.
@@
Having defined our windowed noise schedule $\bm{\bar{\sigma}}({t})=(\bar{\sigma}_{1}({t}),\ldots,\bar{\sigma}_{W}({t}))$, we can formulate the associated probability flow ODE [51], which describes the addition and removal of noise when moving the diffusion ‘time’, ${t}$, forward ($1\rightarrow 0$) and backward ($0\rightarrow 1$), respectively. For a noisy sequence $\bar{\bm{x}}:=\bar{\bm{x}}_{1:W}$, this ODE, as derived in Section C.1, is:
@@
where $\dot{\sigma}_{w}({t})$ denotes the diffusion time derivative of $\bar{\sigma}_{w}({t})$, and $\nabla_{\bar{\bm{x}}}\log p(\bar{\bm{x}};\bm{\bar{\sigma}}({t}))$ is the score function [21] that we will approximate with a denoiser network. This ODE governs $W$ coupled diffusion processes evolving simultaneously within the window. During inference (i.e., the backward process), an ODE solver is used, typically for N discrete steps. Upon completion of one such integration to ${t}=1$, the first snapshot, $\bar{\bm{x}}_{1}(1)$, is fully denoised (its noise level $\bar{\sigma}_{1}(1)=\sigma_{\text{min}}$) and is output as the forecast of ${\bm{y}}_{1}$. Concurrently, the subsequent snapshots $\bar{\bm{x}}_{2:W}(1)$ are partially denoised. Key to the rolling mechanism is the design of the per-snapshot noise schedules $\bar{\sigma}_{w}({t})$, which ensure the noise level of $\bar{\bm{x}}_{w}(1)$ matches the initial noise level of the $(w-1)$-th slot (i.e., $\bar{\sigma}_{w}(1)=\bar{\sigma}_{w-1}(0)$). This specific noise structure at ${t}=1$ enables the window to be advanced efficiently. For the next forecasting step, the sequence $\bar{\bm{x}}_{2:W}(1)$ forms the initial state for the first $W-1$ slots at the next iteration’s ${t}=0$. A new, fully-noised snapshot, $\bm{\epsilon}_{W+1}\sim\mathcal{N}(\mathbf{0},\sigma_{\text{max}}^{2}\mathbf{I}_{D})$, is then appended as the $W$-th element of this new window. The ODE solving process from ${t}=0$ to $1$ is then repeated for this updated window. This iterative sampling procedure is visualized in Fig. 1.

### v3

# Elucidated Rolling Diffusion Models for Probabilistic Forecasting of Complex Dynamics

@@ Abstract
Diffusion models are a powerful tool for probabilistic forecasting, yet most applications in high-dimensional complex systems predict future states individually. This approach struggles to model complex temporal dependencies and fails to explicitly account for the progressive growth of uncertainty inherent to the systems. While rolling diffusion frameworks, which apply increasing noise to forecasts at longer lead times, have been proposed to address this, their integration with state-of-the-art, high-fidelity diffusion techniques remains a significant challenge. We tackle this problem by introducing Elucidated Rolling Diffusion Models (ERDM), the first framework to successfully unify a rolling forecast structure with the principled, performant design of Elucidated Diffusion Models (EDM). To do this, we adapt the core EDM components–its noise schedule, network preconditioning, and Heun sampler–to the rolling forecast setting. The success of this integration is driven by three key contributions: $(i)$ a novel loss weighting scheme that focuses model capacity on the mid-range forecast horizons where determinism gives way to stochasticity; $(ii)$ an efficient initialization strategy using a pre-trained EDM for the initial window; and $(iii)$ a bespoke hybrid sequence architecture for robust spatiotemporal feature extraction under progressive denoising. On 2D Navier–Stokes simulations and ERA5 global weather forecasting at $1.5^{\circ}$ resolution, ERDM consistently outperforms key diffusion-based baselines, including conditional autoregressive EDM. ERDM offers a flexible and powerful general framework for tackling diffusion-based dynamics forecasting problems where modeling uncertainty propagation is paramount. [^1] Code is available at: https://github.com/NVlabs/ERDM
@@ 1 Introduction
Probabilistic forecasting of complex dynamical systems is essential for realistically assessing future states (“snapshots”) and their inherent uncertainties [bauer2015thequiet, clark2001ecology]. For example, medium-range weather forecasting ($\leq 15$ days) grapples with the inherent chaotic nature of the atmosphere and high sensitivities to the initial conditions. Ensembles of numerical weather prediction (NWP) models are therefore standard practice for estimating forecast uncertainty and the likelihood of high-impact events [palmer2002economic, bevacqua2023smiles]. However, each ensemble run of a flagship system such as IFS ENS [ifs-manual-cy46r1-ens] is computationally expensive, motivating data-driven alternatives [bouallegue2024rise].
@@
Diffusion models have shown notable success in generative modeling, particularly for high-dimensional data such as images and videos [sohldickstein2015deepunsupervised, dhariwal2021diffusion]. The EDM framework [karras2022edm], for example, enables stable high-fidelity sample generation due to its principled design that generalizes and improves on the foundational denoising diffusion probabilistic model (DDPM) [ho2020ddpm] and related diffusion paradigms [song2021ddim, song2021scoreSDE]. This principled approach makes them especially promising for modeling time-evolving dynamical systems. In this domain, the ability to generate sharp, diverse ensembles is crucial for robust uncertainty quantification, and producing high-fidelity, physically consistent samples is a key advantage over deterministic models that often yield blurry, averaged-out predictions.
@@
Autoregressive conditional diffusion models have been successfully applied to fluid dynamics [kohl2023benchmarking] and weather [price2023gencast]. These approaches often use limited temporal context and predict future snapshots one by one. To explicitly model the progressively increasing uncertainty common to sequence generation tasks, Rolling Sequence Diffusion Models (RSDM) [ruhe2024rolling, wu2023ardiffusion] employ snapshot-dependent noise schedules that escalate noise for distant future snapshots. However, existing RSDM’s are derived from DDPM, rather than leveraging improved practical design choices, such as network preconditioning and second-order sampling, that were key to EDM’s success.
Figure 1: ERDM sampling with a window, highlighted in bold purple, of size $W=4$. Top row: ERDM starts at diffusion time t=0{t}=0 with snapshots, ${\bm{x}}_{1},\dots,{\bm{x}}_{W}$, corrupted by progressively larger noise levels, $\bar{\sigma}_{1}(t)<\dots<\bar{\sigma}_{W}(t)=\sigma_{\text{max}}$. Middle row: After $N=2$ joint denoising steps, the sequence reaches lower noise levels at t=1{t}=1 such that $\bar{\sigma}_{1}(1)=\sigma_{\text{min}}$ and $\bar{\sigma}_{w}(1)=\bar{\sigma}_{w-1}(0)$ for $w>1$, as illustrated in the right-hand panel. The now fully denoised first snapshot, ${\bm{x}}_{1}$, is returned. Bottom row: The rest of the sequence is shifted one slot to the right, and a fresh pure-noise snapshot is appended to the new window. The cycle then repeats.
@@
Our work addresses these limitations by building upon the successful EDM framework [karras2022edm]. We adapt and extend EDM’s core design elements—its noise schedule formulation, denoiser network parameterization, network preconditioning, loss weighting, and sampling algorithms—to the specific demands of RSDM-like sequence modeling. In developing ERDM, we identified and addressed key design choices for progressive noise schedules and temporal loss weighting within the EDM paradigm, aspects that received limited attention in prior work. Our work thus provides novel solutions for these crucial components. All in all, our key contributions are summarized as follows:
@@ 2 Related Works
Diffusion models, initially transformative in image and audio generation, are increasingly being adapted for spatiotemporal modeling [yang2023diffusion, harvey2022flexiblevideos, gao2023prediff, cachay2023dyffusion]. Latent-diffusion variants [vahdat2021ldm, rombach2022stablediffusion] provide the computational advantages integral to state-of-the-art video generators [ho2022imagenvideo, blattmann2023stablevideodiffusion]. Recent work refines the diffusion process itself to strengthen temporal structure and boost efficiency [cachay2023dyffusion, lippe2023pderefiner]. Rolling sequence diffusion models (RSDM) advance this idea by assigning progressively larger noise levels to later timesteps, mirroring the growth of predictive uncertainty [wu2023ardiffusion, ruhe2024rolling, kim2024fifodiffusion]. We embrace the same intuition but depart from these DDPM-based RSDM’s in two crucial ways. (i) Our approach is derived from the EDM framework, which generalises and improves upon DDPM. (ii) We re-examine often-overlooked design choices—noise schedule, loss weighting, and network architecture—and show that their coordination is critical for forecasting complex dynamics. Progressive schedules like ours form a specific instance of the completely randomized schedules proposed in chen2024diffusionforcing, corresponding to its ‘pyramid sampling’ scheme. However, our ablations indicate that such training-time randomization degrades performance–possibly because it misaligns with the temporally increasing uncertainty inherent to the chaotic systems that we study.
@@
Initial machine learning (ML)-based attempts in this area were limited to deterministic forecasting [pathak2022fourcastnet, bi2023accurate, lam2023learning, nguyen2023climax, nguyen2023stormer, bodnar2024aurora], often differentiated by their underlying neural network architectures [karlbauer2024comparing]. More recently diffusion and flow matching models have shown promise for probabilistic weather forecasting [price2023gencast, couairon2024archesweathergen, andrae2025continuous], and other tasks such as downscaling [mardani2025corrdiff, lopez2025dynamical, chen2023swinrdm, srivastava2024precip], data assimilation [huang2024diffda, rozet2023scorebased, manshausen2025generative], and emulation [li2024genemulation, cachay2024probemulation, pathak2024stormcast]. Latent-variable models offer an alternative probabilistic route [oskarsson2024probabilistic], and physics-ML hybrids such as NeuralGCM show promise but face scalability constraints [kochkov2023neural]. Of particular relevance to our work is GenCast [price2023gencast], which successfully applied the EDM framework, using the graph architecture from [lam2023learning], for next-step probabilistic weather forecasting. Our primary baseline in this paper, EDM, corresponds to the same approach but using our experimental setup (e.g., same U-Net architecture and dataset resolution; see Section˜D.4 for more details). This design ensures a controlled and direct comparison between ERDM and an EDM-based next-step forecasting approach.
@@ 3 Background
In the following, we refer to clean data as $\bm{y}$}, noisy data as $\bm{x}$}, and we use $\bar{\bm{x}}$ to make clear when the data is corrupted according to progressively increasing noise levels. We abbreviate sequences by ${\bm{y}}_{1:W}:={\bm{y}}_{1},\dots,{\bm{y}}_{W}$. Probabilistic forecasting is concerned with predicting a future sequence $\{{\bm{y}}_{\tau}\}_{\tau=1}^{T_{\mathrm{forecast}}}$ given an initial condition ${\bm{y}}_{0}$, where the horizon $T_{\mathrm{forecast}}$ can be arbitrarily large. This requires learning the joint conditional distribution, which can be approximated with, e.g., diffusion models, by $p_{\theta}({\bm{y}}_{1:T_{\mathrm{forecast}}}|{\bm{y}}_{0})$. Directly modeling the high-dimensional joint distribution is often intractable. A common alternative is modeling a smaller window of size $W$ by learning a short-term transition, $p_{\theta}({\bm{y}}_{w+1:w+W}|{\bm{y}}_{w})$ [price2023gencast, harvey2022flexiblevideos]. Long sequences are then generated autoregressively.
@@
Conditional diffusion models can be applied to this task by learning a denoiser network $D_{\theta}({\bm{x}}_{1:W};\sigma,{\bm{y}}_{0})$. The denoiser is trained to predict the target sequence, ${\bm{y}}_{1:W}$, given the initial condition, ${\bm{y}}_{0}$, and data ${\bm{x}}_{1:W}$ corrupted with Gaussian noise of standard deviation $\sigma\in[\sigma_{\text{min}},\sigma_{\text{max}}]$. The maximum noise level needs to be chosen such that $\sigma_{\text{max}}\gg\sigma_{\text{data}}$, where $\sigma_{\text{data}}$ is the standard deviation of the data ${\bm{y}}\sim p_{\rm{data}}$ and ${\bm{y}}\in\mathbb{R}^{D}$. The trained network can then be used jointly with an ODE or SDE solver to iteratively generate the clean data starting from pure noise, $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\sigma_{\text{max}}^{2}\mathbf{I}_{W\times D})$, and ${\bm{y}}_{0}$. See Appendix˜B for an additional background on diffusion. Often, $W=1$ [price2023gencast, kohl2023benchmarking], and the diffusion model becomes a next-step forecasting model. However, the short window size models limited temporal interactions and generating each step requires a full reverse diffusion process, incurring substantial computational cost. Standard sequence diffusion models with $W>1$ typically apply noise uniformly across the prediction window, failing to capture the increasing uncertainty inherent to future predictions.
@@
Rolling Sequence Diffusion Models (RSDMs) [ruhe2024rolling, chen2024diffusionforcing] address prior limitations by explicitly incorporating progressive forecasting uncertainty. The core innovation is a progressive noise schedule with $W>1$. Instead of uniform noise, RSDMs use a monotonically increasing schedule of noise standard deviations $0\approx\bar{\sigma}_{1}^{\text{}}({t})<\bar{\sigma}_{2}^{\text{}}({t})<\dots<\bar{\sigma}_{W}^{\text{}}({t})\approx\sigma_{\text{max}}$, where t{t} denotes the global diffusion time. A noisy future state is thus modeled as $\bar{\bm{x}}_{w}\sim\mathcal{N}({\bm{y}}_{w},\bar{\sigma}^{2}_{w}({t})\mathbf{I}_{D})$. This structure reflects higher uncertainty for predictions further into the future. The denoiser $D_{\theta}(\bar{\bm{x}}_{1:W};\bm{\bar{\sigma}}_{1:W}(t))$ is trained analogously to a standard diffusion model to predict the clean sequence, ${\bm{y}}_{1:W}$. In our implementation, we omit direct conditioning of the denoiser on the last known clean state, ${\bm{y}}_{0}$. This is feasible because we initialize the first window using an external forecaster (itself conditioned on ${\bm{y}}_{0}$). The minimally noisy start of the window provides sufficient conditioning to predict subsequent snapshots. When requiring the RSDM to self-initialize its first window, this is not possible [ruhe2024rolling].
@@ 4 Forecasting with Elucidated Rolling Diffusion Models
Our work, ERDM, integrates the rolling diffusion concept with the Elucidated Diffusion Model (EDM) framework [karras2022edm]. While RSDMs introduce the idea of progressive noise, they usually do not follow EDM’s principled design choices for preconditioning, loss weighting, and sampler design, which we hypothesize would similarly improve RSDM performance. ERDM aims to systematically port and adapt these EDM principles to the rolling, progressive noise setting, thereby providing a robust and theoretically grounded approach to sequential data generation.
@@
Figure 2: Noise schedule comparison for a sequence length $W=6$, yielding 6 visualized segments $\bar{\sigma}_{w}(1),\bar{\sigma}_{w}(0)]$. The schedules differ by their curvature parameter $\rho$: one using $\rho=7$ (default EDM) and the other $\rho=-10$ (ERDM). Color gradients illustrate segment progression.
where $t_{w,t}=1-\frac{w-t}{W}$ is the local diffusion time of the snapshot. The 11- is important so that near-future snapshots receive minimal amounts of noise, while far-future snapshots receive high levels of noise. This ensures the correct inductive bias that distant snapshots are more uncertain than near-future ones. The continuous noise schedule is effectively divided into $W$ segments such that the noise levels at snapshot $w$ satisfy $\bar{\sigma}_{w}(t)\in[\bar{\sigma}_{w-1}(0),\bar{\sigma}_{w+1}(1)]$, where $t\in[0,1],\bar{\sigma}_{0}(0):=\sigma_{\text{min}},$ and $\bar{\sigma}_{W+1}(1):=\sigma_{\text{max}}$. Our formulation thus satisfies $\sigma_{\text{min}}=\bar{\sigma}_{1}(1)<\bar{\sigma}_{1}(0)=\bar{\sigma}_{2}(1)<\dots<\bar{\sigma}_{W}(0)=\sigma_{\text{max}}$. Two example noise schedules are illustrated in Fig.˜2. ERDM is trained to see any noise schedule level within the noise segments (color gradients in Fig.˜2) by randomizing $t\in U([0,1))$. During sampling, each $\bar{\sigma}_{w}({t})$ is evolved from $\bar{\sigma}_{w}(0)$ to $\bar{\sigma}_{w}(1)$. We found it important to tune the curvature parameter, $\rho$. Our proposed default value, $\rho=-10$, effectively causes snapshots to be under less noise compared to the default EDM choice, $\rho=7$, which provides more information for accurate denoising. Importantly, these noise schedule parameters $\sigma_{\text{min}},\sigma_{\text{max}},$ and $\rho$ need to be chosen at training time.
@@
Having defined our windowed noise schedule $\bm{\bar{\sigma}}({t})=(\bar{\sigma}_{1}({t}),\ldots,\bar{\sigma}_{W}({t}))$, we can formulate the associated probability flow ODE [song2021scoreSDE], which describes the addition and removal of noise when moving the diffusion ‘time’, t{t}, forward ($1\rightarrow 0$) and backward ($0\rightarrow 1$), respectively. For a noisy sequence $\bar{\bm{x}}:=\bar{\bm{x}}_{1:W}$, this ODE, as derived in Section˜C.1, is:
@@
where $\dot{\sigma}_{w}({t})$ denotes the diffusion time derivative of $\bar{\sigma}_{w}({t})$, and $\nabla_{\bar{\bm{x}}}\log p(\bar{\bm{x}};\bm{\bar{\sigma}}({t}))$ is the score function [hyvarinen05score] that we will approximate with a denoiser network. This ODE governs $W$ coupled diffusion processes evolving simultaneously within the window. During inference (i.e., the backward process), an ODE solver is used, typically for N discrete steps. Upon completion of one such integration to t=1{t}=1, the first snapshot, $\bar{\bm{x}}_{1}(1)$, is fully denoised (its noise level $\bar{\sigma}_{1}(1)=\sigma_{\text{min}}$) and is output as the forecast of ${\bm{y}}_{1}$. Concurrently, the subsequent snapshots $\bar{\bm{x}}_{2:W}(1)$ are partially denoised. Key to the rolling mechanism is the design of the per-snapshot noise schedules $\bar{\sigma}_{w}({t})$, which ensure the noise level of $\bar{\bm{x}}_{w}(1)$ matches the initial noise level of the (w1)(w-1)-th slot (i.e., $\bar{\sigma}_{w}(1)=\bar{\sigma}_{w-1}(0)$). This specific noise structure at t=1{t}=1 enables the window to be advanced efficiently. For the next forecasting step, the sequence $\bar{\bm{x}}_{2:W}(1)$ forms the initial state for the first W1W-1 slots at the next iteration’s t=0{t}=0. A new, fully-noised snapshot, $\bm{\epsilon}_{W+1}\sim\mathcal{N}(\mathbf{0},\sigma_{\text{max}}^{2}\mathbf{I}_{D})$, is then appended as the $W$-th element of this new window. The ODE solving process from t=0{t}=0 to $1$ is then repeated for this updated window. This iterative sampling procedure is visualized in Fig.˜1.